If the standard deviation of a national accounting examination is 32, how large a sample is needed to estimate the true mean score within 5 points with 98% confidence?
1
Expert's answer
2020-10-21T17:17:14-0400

Standard deviation = 32

z(alpha) = 2.33 - table value for 98% confidence

error = 5

n = ( z(alpha) * SD /error)2= ( 2.33 * 32 / 5 )2=222

sample of 222 is needed.